a vendor sells hot dogs and bags of potato chips. a customer buys 4 hot dogs and 5 bags of potato chips for $12.00. Another customer buys 3 hot dogs and 4 bags of potato chips for $9.25. Find the cost of each of them.
Answer by Boreal(15235)

4H+5C=12.00
3H+4C=9.25
multiply top by 3 and the bottom by -4 to eliminate H
12H+15C=36.00
-12H-16C=-37.00
-C=-1
C=$1 per chips
therefore,
H=$1.75 each